EDITORIAL

Terry Goodman

The Editorial pages of this issue will be devoted to recognizing those individuals who serve as referees for MJMS. We deeply appreciate their contribution to this journal and we recognize that their role is of vital importance. By the way, if you are interested in becoming a referee for MJMS, please fill out the Referee Application Form found on page 161.

You will also find that this issue of MJMS contains a five-year Author and Title Index as well as a five-year Problems Index. We hope you will find these to be useful.

Finally, new subscription rates will go into effect with the February 1994 issue. We regret having to raise the rates, but our expenses have increased over the last five years and we simply have to make these adjustments. We appreciate your support and cooperation.

List of Referees

Hadi Y. Alkahby, Dillard University Charles J. Allard, Polo R-VII High School Mohammad K. Azarian, University of Evansville John E. Bernard, University of Texas–Pan American Albert Dixon, College of the Ozarks John J. Edgell, Jr., Southwest Texas State University Vince Edmondson, Central Missouri State University Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Martin Erickson, Northeast Missouri State University Russell Euler, Northwest Missouri State University Joe Flowers, Northeast Missouri State University Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, Northwest Missouri State University Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh Homer Hampton, Central Missouri State University Dennis R. Harmon, Missouri Southern State College Stan Hartzler, Northwest Missouri State University Troy Hicks, University of Missouri–Rolla Larry K. Johnson, Central Missouri State University

List of Referees (cont.)

N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State College Roger Marty, Cleveland State University Wayne McDaniel, University of Missouri–St. Louis Gary McDonald, Northwest Missouri State University Rhonda L. McKee, Central Missouri State University Kandasamy Muthvel, University of Wisconsin–Oshkosh Ollie Nanyes, Bradley University Krzysztof Ostaszewski, University of Louisville Bob Prielipp, University of Wisconsin–Oshkosh Don Redmond, Southern Illinois University Jeffrey A. Rosoff, Gustavus Adolphus College Jau–Shyong Shiue, University of Nevada–Las Vagas Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Alvin Tinsley, Central Missouri State University Bobby Neal Winters, Pittsburg State University